OpenStudy (theedwardsfamily):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (perl):

for which operations, if you operate, do you get a result that is in the set {-1,1} ?

OpenStudy (theedwardsfamily):

will medal and fan!

OpenStudy (perl):

lets try addition: -1 + 1 = 0 , but 0 is not in the set {-1,1}

OpenStudy (theedwardsfamily):

@perl divison

OpenStudy (theedwardsfamily):

?

OpenStudy (theedwardsfamily):

@sleepyjess

OpenStudy (perl):

yes division works

OpenStudy (zzr0ck3r):

what does closed with resect to division mean?

OpenStudy (theedwardsfamily):

and multiplication, right?

OpenStudy (zzr0ck3r):

can you figure that one out?

OpenStudy (theedwardsfamily):

maybe.

OpenStudy (zzr0ck3r):

what does closed mean?

OpenStudy (theedwardsfamily):

a set that includes all the values obtained by application of a given operation to its members. @zzr0ck3r

OpenStudy (zzr0ck3r):

that would be a closed set so for multiplication, is it true that if we multiply any two elements in the set, we get back another element in the set? hint: 1*1=? 1*-1=? -1*1=? -1*-1*=?